# PID Accelerated Temporal Difference Algorithms

### v1

@@ Abstract
Long-horizon tasks, which have a large discount factor, pose a challenge for most conventional reinforcement learning (RL) algorithms. Algorithms such as Value Iteration and Temporal Difference (TD) learning have a slow convergence rate and become inefficient in these tasks. When the transition distributions are given, PID VI was recently introduced to accelerate the convergence of Value Iteration using ideas from control theory. Inspired by this, we introduce PID TD Learning and PID Q-Learning algorithms for the RL setting in which only samples from the environment are available. We give theoretical analysis of their convergence and acceleration compared to their traditional counterparts. We also introduce a method for adapting PID gains in the presence of noise and empirically verify its effectiveness.
@@ 1 Introduction
The Value Iteration (VI) algorithm is one of the primary dynamic programming methods for solving (discounted) Markov Decision Processes (MDP). It is the foundation of many Reinforcement Learning (RL) algorithms such as the Temporal Difference (TD) Learning (Sutton, 1988; Tsitsiklis and Van Roy, 1997), Q-Learning (Watkins, 1989), Approximate Value/Fitted Iteration (Gordon, 1995; Ernst et al., 2005; Munos and Szepesvári, 2008; Tosatto et al., 2017), and DQN (Mnih et al., 2015; Van Hasselt et al., 2016), which can all be seen as sample-based variants of VI. A weakness of the VI algorithm and the RL algorithms built on top of it is their slow convergence in problems with discount factor $\gamma$ close to $1$, which corresponds to the long-horizon problems where the agent aims to maximize its cumulative rewards far in the future. One can show that the error of the value function calculated by VI at iteration $k$ goes to zero with the rate of ${\mathcal{O}}(\gamma^{k})$. The adverse dependence of the convergence rate on $\gamma$ also appears in the error analysis of the downstream temporal difference (Szepesvári, 1997; Even-Dar and Mansour, 2003; Wainwright, 2019) and fitted value iteration algorithms (Munos and Szepesvári, 2008; Farahmand et al., 2010; Chen and Jiang, 2019; Fan et al., 2019). If $\gamma\approx 1$, these algorithms become very slow and inefficient. This work introduces accelerated temporal difference learning algorithms that can mitigate this issue.
Farahmand and Ghavamzadeh (2021) recently suggested that one may view the iterates of VI as a dynamical system itself. This opens up the possibility of using tools from control theory to modify, and perhaps accelerate, the VI’s dynamics. They specifically used the simple class of Proportional-Integral-Derivative (PID) controllers to modify VI, resulting in a new procedure called the PID VI algorithm. They showed that with a careful choice of the controller gains, PID VI can converge significantly faster than the conventional VI. They also introduced a gain adaptation mechanism, a meta-learning procedure, to automatically choose these gains.
@@ 2 Background
Given a set $\Omega$, let $\mathcal{M}(\Omega)$ be the set of probability distributions over $\Omega$, and $\mathcal{B}(\Omega)$ be the set of bounded functions over $\Omega$. We consider a discounted Markov Decision Process (MDP) (Bertsekas and Tsitsiklis, 1996; Szepesvári, 2010; Sutton and Barto, 2018) defined as $({\mathcal{X}},{\mathcal{A}},\mathcal{P},\mathcal{R},\gamma)$ where ${\mathcal{X}}$ is the finite set of $n$ states, ${\mathcal{A}}$ is the finite set of $m$ actions, $\mathcal{P}\colon{\mathcal{X}}\times{\mathcal{A}}\to\mathcal{M}({\mathcal{X}})$ is the transition kernel, $\mathcal{R}:{\mathcal{X}}\times{\mathcal{A}}\to\mathcal{M}([0,1])$ is the reward function, and $\gamma\in[0,1)$ is the discount factor.
A policy $\pi$ is a function $\pi\colon\mathcal{X}\to\mathcal{M}(\mathcal{A})$ representing the distribution over the actions an agent would take from each state. Given a policy $\pi$, the functions $V^{\pi}:{\mathcal{X}}\to\mathbb{R}$ and $Q^{\pi}\colon{\mathcal{X}}\times{\mathcal{A}}\to\mathbb{R}$ are the corresponding (state-)value function and action-value functions defined as the expected discounted return when following $\pi$ starting at a certain state or state-action pair. We also let ${\mathcal{P}}^{\pi}\colon{\mathcal{X}}\to\mathcal{M}({\mathcal{X}})$ and $\mathcal{R}^{\pi}\colon{\mathcal{X}}\to\mathcal{M}([0,1])$ be the associated transition and reward kernels of policy $\pi$, and $r^{\pi}:{\mathcal{X}}\to[0,1]$ be the expected reward of following $\pi$ at any state.
@@
$(T^{\pi}V)(x)\triangleq r^{\pi}(x)+\gamma\int\mathcal{P}^{\pi}(\text{d}y\mid x)V(y),$ $(\forall x\in{\mathcal{X}})$
@@
The Bellman residuals are functions defined as $\text{BR}^{\pi}V=T^{\pi}V-V$ (for PE) and $\text{BR}^{*}Q={T^{\star}}Q-Q$ (for Control). We write BR to refer to the appropriate choice between $\text{BR}^{\pi}$ and $\text{BR}^{*}$ according to the problem in hand. The value function $V^{\pi}$ is the unique function with $\text{BR}^{\pi}V^{\pi}=0$ and $Q^{*}$ is the unique function with $\text{BR}^{*}Q^{*}=0$.
The iteration $V_{k+1}\leftarrow T^{\pi}V_{k}$ converges to $V^{\pi}$, and the iteration $Q_{k+1}\leftarrow{T^{\star}}Q_{k}$ converges to $Q^{*}$. This is known as the Value Iteration (VI) algorithm. The convergence is due to the $\gamma$-contraction of the Bellman operators with respect to (w.r.t.) the supremum norm, and can be proven using the Banach fixed-point theorem. The result also shows that the convergence rate of VI is $\mathcal{O}(\gamma^{k})$. This can be extremely slow for long horizon tasks with $\gamma$ very close to 1.
@@ 2.1 PID Value Iteration
with $z_{1}$ initialized to a vector of all zeroes. PID VI updates the value function by
@@
The choice of the gains that (maximally) accelerates convergence depends on the MDP and the policy being evaluated. One approach is to place assumptions on the structure of the MDP, and analytically derive the gains that optimize the convergence rate. Farahmand and Ghavamzadeh (2021) provide such a result for the class of reversible Markov chains for the PE problem. Assuming structure on the MDP is not desirable though, so the same work also proposes a gain adaptation algorithm that automatically tunes the controller gains during the fixed point iteration.
@@
The Control case is described similarly by substituting $\text{BR}^{\pi}$ with $\text{BR}^{*}$. The PID VI algorithm (1)–(2) and its gain adaptation procedure (4) depend on the computation of the Bellman residual $\text{BR}^{\pi}V$ or its gradient $\partial\text{BR}^{\pi}V/\partial\kappa_{\cdot}$, both of which require accessing the transition dynamics $\mathcal{P}$. PID VI, like VI, is a dynamic programming/planning algorithm after all. They are not directly applicable to the RL setting, where the agent has access only to samples from a stochastic environment that are obtained online. The goal of the next few sections is to develop RL variants of these algorithms.
@@ 3 PID TD Learning and PID Q-Learning
The only term in $(L_{g}^{\pi}\tilde{V}_{t})(X_{t},\texttt{f})$ that requires estimation is $(\text{BR}^{\pi}V_{t})(X_{t})$, which depends on the transition distributions of the MDP and is not available. We can form an unbiased estimate $\widehat{\text{BR}}_{t}$ of $(\text{BR}^{\pi}V_{t})(X_{t})=(T^{\pi}V)(X_{t})-V(X_{t})$ or $(\text{BR}^{*}Q_{t})(X_{t},A_{t})=({T^{\star}}Q_{t})(X_{t},A_{t})-Q_{t}(X_{t},A_{t})$ by
$\widehat{\text{BR}}_{t}=\begin{cases}R_{t}+\gamma V_{t}(X_{t}^{\prime})-V_{t}(X_{t})&\text{(PE)}\\ R_{t}+\gamma\max_{a^{\prime}\in\mathcal{A}}Q_{t}(X_{t}^{\prime},a^{\prime})-Q_{t}(X_{t},A_{t})&\text{(Control)}\end{cases}.$ (8)
A stochastic approximation procedure can then be used to update the values $V_{t}(X_{t})$, $z_{t}(X_{t})$, $V^{\prime}_{t}(X_{t})$. The procedure would be
@@
We call this procedure the PID TD learning algorithm. For Control, we similarly form estimates $L_{t,\texttt{f}}$ of $(L_{g}^{*}Q_{t})(X_{t},A_{t})$ and obtain
@@
where $\mu_{t}=\mu(N_{t}(X_{t},A_{t}))$. This is the PID Q-Learning algorithm. It is worth mentioning that one can use other forms of learning rates, such as constant or state-count independent or choosing three different learning rates for the three updates in (3) and (3) to achieve better practical results. The formulation in this section is chosen for simplicity and theoretical analysis.
@@ 4.1 Convergence Guarantee
Farahmand and Ghavamzadeh (2021) show that PID VI converges under a wide range of gains for a wide range of environments both analytically and experimentally. We show that this convergence carries over to our sample-based PID TD Learning. We first need to define some notation to express our result. Note that $L_{g}^{\pi}$ is an affine linear operator. Define $A_{g}^{\pi}$ to be its linear component and $b_{g}^{\pi}$ to be the constant component, so that $L_{g}^{\pi}\tilde{V}=A_{g}^{\pi}\tilde{V}+b_{g}^{\pi}$. In particular,
@@
The matrix $A_{g}^{\pi}$ plays a critical role in the behavior of PID VI as well as PID TD Learning. Farahmand and Ghavamzadeh (2021) show that PID VI is convergent for PE if $\rho(A_{g}^{\pi})<1$ where $\rho(M)$ for a square matrix $M$ is its spectral radius, the maximum of the magnitude of the eigenvalues. It turns out the condition on controller gains needed for convergence of PID TD Learning is weaker than the one for PID VI. We provide the following result.
@@
It is shown that if this ODE converges to the stationary point $\tilde{V}^{\pi}$, PID TD Learning will also converge. The condition for the convergence of this linear ODE is that the eigenvalues $\{\lambda^{\prime}_{i}\}$ of $A_{g}^{\pi}-I$ should have negative real parts. Since $\lambda_{i}^{\prime}=\lambda_{i}-1$, we get the condition in Theorem 1. Note that this condition is weaker than $\rho(A_{g}^{\pi})<1$ for PID VI Farahmand and Ghavamzadeh (2021), which is equivalent to $|\lambda_{i}|<1$. In other words, PID TD Learning maybe convergent even if PID VI with the same controller gains $g$ is not.
Obtaining similar results for PID Q-Learning is technically much more challenging. The reason for this difficulty is the fact that, just like PID VI for Control, as the agent’s policy changes, the dynamics of PID Q-Learning changes. Moreover, as opposed to Bellman operator $T^{*}$ that is a contraction, $L_{g}^{*}$ is not always a contraction. Similar to Farahmand and Ghavamzadeh (2021), we leave theoretical analysis of PID Q-Learning to future work and only focus on its empirical study.
@@ 4.2 Acceleration Result
In this section, we provide theoretical insights on how PID TD Learning can show a faster convergence compared to the conventional TD Learning. Our analysis relies on the finite-sample analysis of stochastic approximation methods. Since results for the asynchronous updates are limited, we provide our acceleration results for synchronous updates. Specifically, we provide our analysis for the case that at each iteration $t$, a dataset $\{(x,A_{x},R_{x},X^{\prime}_{x})\}_{x\in{\mathcal{X}}}$ is given, where for each state $x\in{\mathcal{X}}$ it contains the random action $A_{x}\sim\pi(\cdot|x)$, reward $R_{x}\sim\mathcal{R}(x,A_{x})$, and $X^{\prime}_{x}\sim{\mathcal{P}}^{\pi}(\cdot|x,A_{x})$. Then, all values of $V,V^{\prime}$, and $z$ are updated simultaneously in the same manner as (3). Similarly, synchronous TD Learning applies the conventional update on all states using the dataset. Based on the analysis by Chen et al. (2020b), the following theorem provides bounds on the error of both algorithms for the learning rate schedule $\mu(t)=\epsilon/(t+T)$. We focus on the choices of $\epsilon,T$ that achieve the optimal asymptotic rate.
@@
Suppose synchronous TD Learning and synchronous PID TD Learning are run with initial value function $V_{0}$ and learning rate $\mu(t)=\epsilon/(t+T)$ to evaluate policy $\pi$. Let $V^{\mathrm{TD}}_{t},V^{\mathrm{PID}}_{t}$ be the value function obtained by the algorithms at iteration $t$, and $\{c^{\mathrm{TD}}_{i},c^{\mathrm{PID}}_{i}\}$ be constants only dependent on the MDP and controller gains. Assume ${\mathcal{P}}^{\pi}$ is diagonalizable. If $\epsilon>2/(1-\gamma)$ and $T\geq c^{\mathrm{TD}}_{1}\epsilon/(1-\gamma)$, we have
@@
The assumption on diagonalizability of ${\mathcal{P}}^{\pi}$ and $A_{g}^{\pi}$ in Theorem 2 is for the sake of simplicity. In Appendix B, we provide a similar but more general result without this assumption. The upper bounds in Theorem 2 consist of two terms. The first term, which scales with the initial error $\lVert{V_{0}-V^{\pi}}\rVert_{\infty}$ can be interpreted as the optimization error. It is the amount that $V_{t}$ still has to change to reach $V^{\pi}$. The second term can be considered as the statistical error, which is independent of the initial error and exists even if we start from $V_{0}=V^{\pi}$. Due to the conditions $\epsilon>2/(1-\gamma)$ and $\epsilon>2/(1-\rho)$, the statistical error is asymptotically dominant with rate ${\mathcal{O}}(t^{-1})$ compared to ${\mathcal{O}}(t^{-\epsilon(1-\gamma)})$ or ${\mathcal{O}}(t^{-\epsilon(1-\gamma)})$ of the optimization error. Note that a larger $\epsilon$ accelerates the rate of optimization error, but together with larger $T$ (due to the condition on $T$) slows the statistical error. For example, it takes $T$ steps for the statistical error to become half of its initial value. For simplicity of discussion, we consider $\epsilon$ and $T$ fixed.
The difference between the two algorithms is in the rate that the optimization error goes to zero. This term for TD Learning is ${\mathcal{O}}(t^{-\epsilon(1-\gamma)})$ and for PID TD Learning is ${\mathcal{O}}(t^{-\epsilon(1-\rho)})$. When $\kappa_{p}=1$ and $\kappa_{I}=\kappa_{d}=\beta=0$, we have $\rho=\gamma$, and these two rates match. With a better choice of gains, one can have $\rho<\gamma$ (Farahmand and Ghavamzadeh, 2021) and achieve a faster rate for the optimization error. Even though this term is not asymptotically dominant, we show that its speed-up can be significant in the early stages of training, especially when the policy’s behavior has low stochasticity. To show this, we first need to introduce the following definition.

### v2

@@ Abstract
Long-horizon tasks, which have a large discount factor, pose a challenge for most conventional reinforcement learning (RL) algorithms. Algorithms such as Value Iteration and Temporal Difference (TD) learning have a slow convergence rate and become inefficient in these tasks. When the transition distributions are given, PID VI was recently introduced to accelerate the convergence of Value Iteration using ideas from control theory. Inspired by this, we introduce PID TD Learning and PID Q-Learning algorithms for the RL setting, in which only samples from the environment are available. We give a theoretical analysis of the convergence of PID TD Learning and its acceleration compared to the conventional TD Learning. We also introduce a method for adapting PID gains in the presence of noise and empirically verify its effectiveness.
@@ 1 Introduction
The Value Iteration (VI) algorithm is one of the primary dynamic programming methods for solving (discounted) Markov Decision Processes (MDP). It is the foundation of many Reinforcement Learning (RL) algorithms such as the Temporal Difference (TD) Learning (Sutton, 1988; Tsitsiklis and Van Roy, 1997), Q-Learning (Watkins, 1989), Approximate/Fitted Value Iteration (Gordon, 1995; Ernst et al., 2005; Munos and Szepesvári, 2008; Tosatto et al., 2017), and DQN (Mnih et al., 2015; Van Hasselt et al., 2016), which can all be seen as sample-based variants of VI. A weakness of the VI algorithm and the RL algorithms built on top of it is their slow convergence in problems with discount factor $\gamma$ close to $1$, which corresponds to the long-horizon problems where the agent aims to maximize its cumulative rewards far in the future. One can show that the error of the value function calculated by VI at iteration $k$ goes to zero with the slow rate of ${\mathcal{O}}(\gamma^{k})$. The slow convergence rate when $\gamma\approx 1$ also appears in the error analysis of the downstream temporal difference (Szepesvári, 1997; Even-Dar and Mansour, 2003; Wainwright, 2019) and fitted value iteration algorithms (Munos and Szepesvári, 2008; Farahmand et al., 2010; Chen and Jiang, 2019; Fan et al., 2019). If $\gamma\approx 1$, these algorithms become very slow and inefficient. This work introduces accelerated temporal difference learning algorithms that can mitigate this issue.
Farahmand and Ghavamzadeh (2021) recently suggested that one may view the iterates of VI as a dynamical system. This opens up the possibility of using tools from control theory to modify, and perhaps accelerate, the VI’s dynamics. They specifically used the simple class of Proportional-Integral-Derivative (PID) controllers to modify VI, resulting in a new procedure called the PID VI algorithm. They showed that with a careful choice of the controller gains, PID VI can converge significantly faster than the conventional VI. They also introduced a gain adaptation mechanism, a meta-learning procedure, to automatically choose these gains.
@@ 2 Background
Given a set $\Omega$, let $\mathcal{M}(\Omega)$ be the set of probability distributions over $\Omega$, and $\mathcal{B}(\Omega)$ be the set of bounded functions over $\Omega$. We consider a discounted MDP (Bertsekas and Tsitsiklis, 1996; Szepesvári, 2010; Sutton and Barto, 2018) defined as $({\mathcal{X}},{\mathcal{A}},\mathcal{P},\mathcal{R},\gamma)$ where ${\mathcal{X}}$ is the finite set of $n$ states, ${\mathcal{A}}$ is the finite set of $m$ actions, $\mathcal{P}\colon{\mathcal{X}}\times{\mathcal{A}}\to\mathcal{M}({\mathcal{X}})$ is the transition kernel, $\mathcal{R}:{\mathcal{X}}\times{\mathcal{A}}\to\mathcal{M}([0,1])$ is the reward function, and $\gamma\in[0,1)$ is the discount factor.
A policy $\pi$ is a function $\pi\colon\mathcal{X}\to\mathcal{M}(\mathcal{A})$ representing the distribution over the actions an agent would take from each state. Given a policy $\pi$, the functions $V^{\pi}:{\mathcal{X}}\to\mathbb{R}$ and $Q^{\pi}\colon{\mathcal{X}}\times{\mathcal{A}}\to\mathbb{R}$ are the corresponding (state-)value and action-value functions defined as the expected discounted return when following $\pi$ starting at a certain state or state-action pair. We also let ${\mathcal{P}}^{\pi}\colon{\mathcal{X}}\to\mathcal{M}({\mathcal{X}})$ and $\mathcal{R}^{\pi}\colon{\mathcal{X}}\to\mathcal{M}([0,1])$ be the associated transition and reward kernels of policy $\pi$, and $r^{\pi}:{\mathcal{X}}\to[0,1]$ be the expected reward of following $\pi$ at any state.
@@
$(T^{\pi}V)(x)\triangleq r^{\pi}(x)+\gamma\int\mathcal{P}^{\pi}(\text{d}y\mid x)V(y),$ $(\forall x\in{\mathcal{X}}),$
@@
The Bellman residual operators are defined as $\text{BR}^{\pi}V\triangleq T^{\pi}V-V$ (for PE) and $\text{BR}^{*}Q\triangleq{T^{\star}}Q-Q$ (for Control). The value function $V^{\pi}$ is the unique function with $\text{BR}^{\pi}V^{\pi}=0$ and $Q^{*}$ is the unique function with $\text{BR}^{*}Q^{*}=0$.
The iteration $V_{k+1}\leftarrow T^{\pi}V_{k}$ converges to $V^{\pi}$, and the iteration $Q_{k+1}\leftarrow{T^{\star}}Q_{k}$ converges to $Q^{*}$. This is known as the VI algorithm. The convergence is due to the $\gamma$-contraction of the Bellman operators with respect to the supremum norm, and can be proven using the Banach fixed-point theorem. The result also shows that the convergence rate of VI is $\mathcal{O}(\gamma^{k})$. This can be extremely slow for long horizon tasks with $\gamma$ very close to 1.
@@ 2.1 PID Value Iteration
with $\alpha,\beta\in\mathbb{R}$ and $z_{1}$ initialized to a vector of all zeroes. PID VI updates the value function by
@@
The choice of the gains that (maximally) accelerates convergence depends on the MDP and the policy being evaluated. One approach is to place assumptions on the structure of the MDP, and analytically derive the gains that optimize the convergence rate. Farahmand and Ghavamzadeh (2021) provide such a result for PE in the class of reversible Markov chains. Assuming structure on the MDP is not desirable though, so the same work also proposes a gain adaptation algorithm that automatically tunes the controller gains during the fixed point iteration.
@@
The Control case is described similarly by substituting $\text{BR}^{\pi}$ with $\text{BR}^{*}$. The PID VI algorithm (1)–(2) and its gain adaptation procedure (4) depend on the computation of the Bellman residual $\text{BR}^{\pi}V$ or its gradient $\partial\text{BR}^{\pi}V/\partial\kappa_{\cdot}$, both of which require accessing the transition dynamics $\mathcal{P}$. PID VI, like VI, is a dynamic programming/planning algorithm after all. They are not directly applicable to the RL setting, where the agent has access only to samples from the environment that are obtained online. The goal of the next few sections is to develop RL variants of these algorithms.
@@ 3 PID TD Learning and PID Q-Learning
The only term in $(L_{g}^{\pi}\tilde{V}_{t})(X_{t},\texttt{f})$ that requires estimation is $(\text{BR}^{\pi}V_{t})(X_{t})$, which depends on the transition distributions of the MDP that is not available. We can form an unbiased estimate $\widehat{\text{BR}}_{t}$ of $(\text{BR}^{\pi}V_{t})(X_{t})=(T^{\pi}V)(X_{t})-V(X_{t})$ or $(\text{BR}^{*}Q_{t})(X_{t},A_{t})=({T^{\star}}Q_{t})(X_{t},A_{t})-Q_{t}(X_{t},A_{t})$ by
$\widehat{\text{BR}}_{t}=\begin{cases}R_{t}+\gamma V_{t}(X_{t}^{\prime})-V_{t}(X_{t})&\text{(PE)},\\ R_{t}+\gamma\max_{a^{\prime}\in\mathcal{A}}Q_{t}(X_{t}^{\prime},a^{\prime})-Q_{t}(X_{t},A_{t})&\text{(Control)}.\end{cases}$ (8)
A stochastic approximation procedure can then be used to update the values $V_{t}(X_{t})$, $z_{t}(X_{t})$, $V^{\prime}_{t}(X_{t})$ according to (7). The procedure would be
@@
We call this procedure the PID TD learning algorithm. For Control, we similarly form estimates $\hat{L}_{t,\texttt{f}}$ of $(L_{g}^{*}Q_{t})(X_{t},A_{t},\texttt{f})$ and obtain
@@
where $\mu_{t}=\mu(N_{t}(X_{t},A_{t}))$. This is the PID Q-Learning algorithm. It is worth mentioning that one can use other forms of learning rates to achieve better practical results. For example, we can choose constant or state-count independent learning rates, or use three different learning rates for the three updates in (3) and (3). The formulation in this section is chosen for simplicity and the theoretical analysis.
@@ 4.1 Convergence Guarantee
Farahmand and Ghavamzadeh (2021) show that PID VI converges under a wide range of gains for a wide range of environments both analytically and experimentally. We show that this convergence carries over to our sample-based PID TD Learning. We first need to define some notations to express our result. Note that $L_{g}^{\pi}$ is an affine linear operator. Define $A_{g}^{\pi}$ to be its linear component and $b_{g}^{\pi}$ to be the constant component, so that $L_{g}^{\pi}\tilde{V}=A_{g}^{\pi}\tilde{V}+b_{g}^{\pi}$. In particular,
@@
The matrix $A_{g}^{\pi}$ plays a critical role in the behavior of PID VI as well as PID TD Learning. Farahmand and Ghavamzadeh (2021) show that PID VI is convergent for PE if $\rho(A_{g}^{\pi})<1$ where $\rho(M)$ for a square matrix $M$ is its spectral radius, the maximum of the magnitude of the eigenvalues. It turns out the condition on the controller gains needed for the convergence of PID TD Learning is weaker than the one for PID VI. We provide the following result.
@@
It is shown that if this ODE converges to the stationary point $\tilde{V}^{\pi}$, PID TD Learning will also converge. The condition for the convergence of this linear ODE is that the eigenvalues $\{\lambda^{\prime}_{i}\}$ of $A_{g}^{\pi}-I$ should have negative real parts. Since $\lambda_{i}^{\prime}=\lambda_{i}-1$, we get the condition in Theorem 1. Note that this condition is weaker than $\rho(A_{g}^{\pi})<1$ for PID VI (Farahmand and Ghavamzadeh, 2021), which is equivalent to $|\lambda_{i}|<1$. In other words, PID TD Learning may be convergent even if PID VI with the same controller gains $g$ is not.
Obtaining similar results for PID Q-Learning is technically much more challenging. The reason for this difficulty is the fact that, just like PID VI for Control, as the agent’s policy changes, the dynamics of PID Q-Learning changes. Similar to Farahmand and Ghavamzadeh (2021), we leave theoretical analysis of PID Q-Learning to future work and only focus on its empirical study.
@@ 4.2 Acceleration Result
In this section, we provide theoretical insights on how PID TD Learning can show a faster convergence compared to the conventional TD Learning. Our analysis relies on the finite-sample analysis of stochastic approximation methods. Since results for the asynchronous updates are limited, we provide our acceleration results for synchronous updates. Specifically, we provide our analysis for the case that at each iteration $t$, a dataset $\mathcal{D}_{t}=\{(x,A_{x,t},R_{x,t},X^{\prime}_{x,t})\}_{x\in{\mathcal{X}}}$ is given, where for each state $x\in{\mathcal{X}}$ it contains the random action $A_{x,t}\sim\pi(\cdot|x)$, reward $R_{x,t}\sim\mathcal{R}(x,A_{x,t})$, and $X^{\prime}_{x,t}\sim{\mathcal{P}}^{\pi}(\cdot|x,A_{x,t})$. Then, all values of $V,V^{\prime}$, and $z$ are updated simultaneously in the same manner as (3). Similarly, synchronous TD Learning applies the conventional update on all states using the dataset. Based on the analysis by Chen et al. (2020), the following theorem provides bounds on the error of both algorithms for the learning rate schedule $\mu(t)=\epsilon/(t+T)$. We focus on the choices of $\epsilon,T$ that achieve the optimal asymptotic rate.
@@
Suppose synchronous TD Learning and synchronous PID TD Learning are run with initial value function $V_{0}$ and learning rate $\mu(t)=\epsilon/(t+T)$ to evaluate policy $\pi$. Let $V^{\mathrm{TD}}_{t}$ and $V^{\mathrm{PID}}_{t}$ be the value functions obtained by the algorithms at iteration $t$, and $\{c^{\mathrm{TD}}_{i},c^{\mathrm{PID}}_{i}\}$ be constants only dependent on the MDP and controller gains. Assume ${\mathcal{P}}^{\pi}$ is diagonalizable. If $\epsilon>2/(1-\gamma)$ and $T\geq c^{\mathrm{TD}}_{1}\epsilon/(1-\gamma)$, we have
@@
The assumption on diagonalizability of ${\mathcal{P}}^{\pi}$ and $A_{g}^{\pi}$ in Theorem 2 is for the sake of simplicity. In Appendix B, we provide a similar but more general result without this assumption. The upper bounds in Theorem 2 consist of two terms. The first term, which scales with the initial error $\lVert{V_{0}-V^{\pi}}\rVert_{\infty}$ can be interpreted as the optimization error. It is the amount that $V_{t}$ still has to change to reach $V^{\pi}$. The second term can be considered as the statistical error, which is independent of the initial error and exists even if we start from $V_{0}=V^{\pi}$. Due to the conditions $\epsilon>2/(1-\gamma)$ and $\epsilon>2/(1-\rho)$, the statistical error is asymptotically dominant with rate ${\mathcal{O}}(t^{-1})$ compared to ${\mathcal{O}}(t^{-\epsilon(1-\gamma)})$ or ${\mathcal{O}}(t^{-\epsilon(1-\gamma)})$ of the optimization error. Note that a larger $\epsilon$ accelerates the rate of optimization error, but together with larger $T$ (due to the condition on $T$) slows the convergence of the statistical error to zero. For example, it takes $T$ steps for the statistical error to become half of its initial value. For simplicity of discussion, we consider $\epsilon$ and $T$ fixed.
The difference between the two algorithms is in the rate that the optimization error goes to zero. This term for TD Learning is ${\mathcal{O}}(t^{-\epsilon(1-\gamma)})$ and for PID TD Learning is ${\mathcal{O}}(t^{-\epsilon(1-\rho)})$. When $\kappa_{p}=1$ and $\kappa_{I}=\kappa_{d}=\alpha=0$, we have $\rho=\gamma$, and these two rates match. With a better choice of gains, one can have $\rho<\gamma$ (Farahmand and Ghavamzadeh, 2021) and achieve a faster rate for the optimization error. Even though this term is not asymptotically dominant, we show that its speed-up can be significant in the early stages of training, especially when the policy’s behavior has low stochasticity. To show this, we first need to introduce the following definition.